# A Homothetic Gauge Theory

### v1

@@ 4.2 Generalized de Rham complex and Hodge decomposition
$\bar{d}:=\begin{pmatrix}d&0\\ 0&d\end{pmatrix},\quad\bar{\delta}:=\begin{pmatrix}\delta&0\\ 0&\delta\end{pmatrix},\quad\bar{\Delta}:=\begin{pmatrix}\Delta&0\\ 0&\Delta\end{pmatrix},\qquad\ \bar{\star}:=\begin{pmatrix}\star&0\\ 0&\star\end{pmatrix},\quad$ (26)
@@
    With the Weyl-rescaled metric $\tilde{\eta}=e^{-2\lambda(x)}\eta$, $\hat{\delta}$ is the adjoint of $\hat{d}$, that is, $\langle\hat{d}\hat{\alpha},\hat{\beta}\rangle_{\tilde{\eta}}=\langle\hat{\alpha},\hat{\delta}\hat{\beta}\rangle_{\tilde{\eta}}$.
@@
Geometric and Physical Interpretations. The generalized operators $\hat{d},\hat{\delta},\hat{\Delta}$, $\hat{\star}$ are not arbitrary: they arise naturally from the demand that the differential structure be compatible with the homothetic transformations introduced earlier. Specifically, the exterior derivative and codifferential must transform covariantly under the homothetic action $\Lambda$, preserving the structure of the deformed forms $\hat{\phi}=\Lambda\bar{\phi}$. This compatibility is encoded in the identities (31) which imply that homothetic transformations commute with the differential operators, effectively rendering the homothetic structure gauge covariant.
Physically, this suggests that the transformation group $\mathcal{H}$ plays a role analogous to a gauge group, albeit non-compact and infinite-dimensional. The field $\lambda(x)$ behaves like a scalar gauge parameter—more precisely, a logarithmic scale factor—whose differential $d\lambda$ enters the transformation of both $d$ and $\delta$ in a manner reminiscent of a connection 1-form in Weyl geometry. The antisymmetric contributions involving $d\lambda\wedge$ and $*d\lambda\wedge*$ indicate that this is a kind of abelian connection with scale-coupled torsion.
@@
00{0}${\hat{\Omega}^{0}}$${\hat{\Omega}^{1}}$${\hat{\Omega}^{2}}$${\hat{\Omega}^{3}}$${\hat{\Omega}^{4}}$00{0}$\scriptstyle{i}$$\scriptstyle{\hat{\Delta}_{0}}$$\scriptstyle{\hat{\delta}_{0}}$$\scriptstyle{\hat{d}_{0}}$$\scriptstyle{\hat{\Delta}_{1}}$$\scriptstyle{\hat{\delta}_{1}}$$\scriptstyle{\hat{d}_{1}}$$\scriptstyle{\hat{\Delta}_{2}}$$\scriptstyle{\hat{\delta}_{2}}$$\scriptstyle{\hat{d}_{2}}$$\scriptstyle{\hat{\Delta}_{3}}$$\scriptstyle{\hat{\delta}_{3}}$$\scriptstyle{\hat{d}_{3}}$$\scriptstyle{\hat{\Delta}_{4}}$$\scriptstyle{\hat{\delta}_{4}}$$\scriptstyle{\hat{d}_{4}}$$\scriptstyle{i}$
Figure 2: Diagrammatic representation of the generalized de Rham complex with homothetic differential forms $\hat{\Omega}^{k}$, showing both the generalized exterior derivative $\hat{d}$ and codifferential $\hat{\delta}$. At each degree $k$, the kernel of the generalized Laplacian $\hat{\Delta}$ identifies the harmonic representatives $\mathcal{H}^{k}$ in the generalized Hodge decomposition $\mathcal{H}^{k}:=\ker\hat{\Delta}^{k}$ (${\mathcal{H}}^{k}$ not to be confused with the homothety group $\mathcal{H}$).
Cohomological Consequences. Importantly, the nilpotency of $\hat{d}$ allows us to define a generalized de Rham cohomology:
@@ 5 Homothetic Electromagnetism
The above homothetic de Rham complex can be constructed for the electromagnetism, as a $U(1)$ gauge theory defined on the Minkowski spacetime manifold. It means embedding any $k$-form $\alpha^{k}$, where $0\leq k\leq 4$ in $\Omega^{k}(M)\oplus\Omega^{k}(M)$, where $\mathcal{H}$ can act on, making $\hat{\Omega}^{k}$. However, in order for the homothetic transformations to preserve the symmetries of the electromagnetism, they must act directly on the connections of the principal bundle, that is, the gauge potential fields $A=A_{\mu}dx^{\mu}$ (as the $\mathfrak{u}(1)$-valued 1-forms on $M$). Such a setup results in Fig. 3, in which we defined:
@@ 5.1 Sourceless Maxwell Equations
Since $\hat{F}=\hat{d}\hat{A}$, the gauge transformation $\hat{A}\mapsto\hat{A}+\hat{d}\phi$ leaves $\hat{F}$ invariant, in which:
$\hat{\phi}=\begin{pmatrix}\tilde{\phi}\\ \phi_{d}\end{pmatrix}.$ (39)
Substituting into the homothetic Maxwell equations (37)-(38) yields:
@@
As one can see, some additional terms have appeared which represent the interaction of the electromagnetic fields with the offset fields. Apparently, these additional terms disappear, and the classical Maxwell equations retrieve for $d\lambda\equiv 0$.
00{0}${\Omega^{0}}$${\Omega^{1}}$${\Omega^{2}}$${\Omega^{3}}$${\Omega^{4}}$00{0}00{0}${\hat{\Omega}^{0}}$${\hat{\Omega}^{1}}$${\hat{\Omega}^{2}}$${\hat{\Omega}^{3}}$${\hat{\Omega}^{4}}$00{0}$\scriptstyle{i}$$\scriptstyle{\definecolor[named]{.}{rgb}{.5,.5,.5}\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(\mathcal{H}{,}\delta A_{d})}$$\scriptstyle{\Delta_{0}}$$\scriptstyle{\delta_{0}}$$\scriptstyle{d_{0}}$$\scriptstyle{\definecolor[named]{.}{rgb}{1,0,0}\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(\mathcal{H}{,}A_{d})}$$\scriptstyle{\Delta_{1}}$$\scriptstyle{\delta_{1}}$$\scriptstyle{d_{1}}$$\scriptstyle{\definecolor[named]{.}{rgb}{.5,.5,.5}\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(\mathcal{H}{,}dA_{d})}$$\scriptstyle{\Delta_{2}}$$\scriptstyle{\delta_{2}}$$\scriptstyle{d_{2}}$$\scriptstyle{\definecolor[named]{.}{rgb}{.5,.5,.5}\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(\mathcal{H}{,}0)}$$\scriptstyle{\Delta_{3}}$$\scriptstyle{\delta_{3}}$$\scriptstyle{d_{3}}$$\scriptstyle{\definecolor[named]{.}{rgb}{.5,.5,.5}\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(\mathcal{H}{,}0)}$$\scriptstyle{\Delta_{4}}$$\scriptstyle{\delta_{4}}$$\scriptstyle{d_{4}}$$\scriptstyle{i}$$\scriptstyle{i}$$\scriptstyle{\hat{\Delta}_{0}}$$\scriptstyle{\hat{\delta}_{0}}$$\scriptstyle{\hat{d}_{0}}$$\scriptstyle{\hat{\Delta}_{1}}$$\scriptstyle{\hat{\delta}_{1}}$$\scriptstyle{\hat{d}_{1}}$$\scriptstyle{\hat{\Delta}_{2}}$$\scriptstyle{\hat{\delta}_{2}}$$\scriptstyle{\hat{d}_{2}}$$\scriptstyle{\hat{\Delta}_{3}}$$\scriptstyle{\hat{\delta}_{3}}$$\scriptstyle{\hat{d}_{3}}$$\scriptstyle{\hat{\Delta}_{4}}$$\scriptstyle{\hat{\delta}_{4}}$$\scriptstyle{\hat{d}_{4}}$$\scriptstyle{i}$
Figure 3: Homothetic de Rham complex for the classical electromagnetism (the lower part), and its relation with the classical de Rham complex (the upper part). The homothetic transformation genuinely occurs on the connection 1-forms, and it induces other transformations on the other $k\neq 1$ differential forms.
By translation of the above equations on the tangent bundle, the generalized Maxwell equations are obtained in the more conventional vector calculus form:
@@
[…]
@@
[…]
And since both the four-velocity vector $u^{\mu}$ and the electromagnetic field $F^{\mu\nu}$ are modified, the Lorentz force $f^{\mu}_{L}$ changes to:
$\tilde{f}^{\mu}_{L}=e^{-\lambda(x)}q\tilde{F}^{\mu\nu}u_{\nu},$ (46)
showing a modification in the Lorentz force.
This formulation offers a homothetically covariant generalization of electromagnetism. A distinguishing feature of this system is that it is sourceless. In fact, the emergence of penalty terms, with geometric origins, enable us to derive the dynamics of electromagnetic fields without the need for source terms. This point will be discussed further in §$\S$ 6.4.
6 The Homothetic Conservation, Wave, and Laplace Equations
We now turn to the derivation of the dynamical equations that arise from homothetically transformed electromagnetic potentials and scalar fields. These include a conservation law, a wave equation, and a Laplace equation defined on the homothetically transformed geometry.
@@ 6.1 Homothetic Harmonicity of the Gauge Potential
6.2 Conservation Equation
From the generalized Lorenz gauge (47), with:
@@ 6.2 Conservation Equation
This yields a generalized conservation law:
@@
In spacetime index notation, the conservation equation becomes:
@@ 6.3 Wave and Laplace Equations
6.4 About the Offset Fields $\phi_{d}$, $A_{d}$ and $F_{d}$
The remaining parameter to be determined is the offset potential $A_{d}$. Clearly, the source-free electromagnetic field is homothetically invariant, just as it is conformally invariant. That is,
$\boxed{A_{d}=A.}$ (50)
Thus, $A_{d}$ represents the asymptotic, “frozen” gauge potential in regions where causal propagation terminates. This implies that $A_{d}$ should be intrinsically connected to the initial physical potential $A$ itself, rather than being an independent quantity.
The consistency of the homothetic invariance of $A$ can be verified using the homothetic conservation law derived from the Lorenz gauge condition $\hat{\delta}\hat{A}=0$, as stated in Eq. (48):
$\delta\tilde{A}-\star(d\lambda)\wedge\star(\tilde{A}-A_{d})=0.$ (51)
Substituting $\tilde{A}=A$ and $A_{d}=A$ into this equation gives:
$\delta A-\star(d\lambda)\wedge\star(A-A)=0,$ (52)
which simplifies to:
$\delta A=0.$ (53)
This is precisely the standard Lorenz gauge condition for the original potential $A$. Therefore, the identification $A_{d}=A$ is entirely consistent with the homothetic gauge structure, provided that the initial potential $A$ is chosen to satisfy the Lorenz gauge. This choice is not merely a mathematical convenience; rather, it elevates the framework by establishing the physical, Lorenz-gauged potential $A$ as the fundamental invariant quantity that persists across all homothetic scales.
Note that the homothetic invariance of the gauge potentials naturally extends—via the $d$ and $\delta$ operators—to the electromagnetic fields and scalar potentials:
$\boxed{F_{d}=F,\qquad\phi_{d}=\phi.}$ (54)
In §$\S$ 5.1, we noted that the generalized Maxwell equations (40)–(41) or (42)–(45) do not require source terms. We now observe that the offset fields are, in fact, the (frozen) conventional electromagnetic fields. Perhaps this self-contained nature is the most distinctive feature of the present theory, setting it apart from the classical Maxwell equations, which rely on externally defined source terms.
7 Cauchy, Dirichlet and Neumann Boundary Conditions for the Laplace Equation

### v2

@@ 4.2 Generalized de Rham complex and Hodge decomposition
$\bar{d}:=\begin{pmatrix}d&0\\ 0&d\end{pmatrix},\quad\bar{\delta}:=\begin{pmatrix}\delta&0\\ 0&\delta\end{pmatrix},\quad\bar{\Delta}:=\begin{pmatrix}\Delta&0\\ 0&\Delta\end{pmatrix},\qquad\ \bar{\star}:=\begin{pmatrix}\star&0\\ 0&\star\end{pmatrix}$$\quad$ (24)
@@
    With the conformally scaled metric $\tilde{\eta}=e^{-2\lambda(x)}\eta$, $\hat{\delta}$ is the adjoint of $\hat{d}$  [^2] It should be noted that without the conformal scaling $\eta\to\tilde{\eta}$, $\hat{\delta}$ is not the formal adjoint of $\hat{d}$. It may be seen as an evidence of internal consistency of the theory., that is:
    $\langle\hat{d}\hat{\alpha},\hat{\beta}\rangle_{\tilde{\eta}}=\langle\hat{\alpha},\hat{\delta}\hat{\beta}\rangle_{\tilde{\eta}}$ (30)
@@
Geometric and Physical Interpretations
The generalized operators $\hat{d},\hat{\delta},\hat{\Delta}$, $\hat{\star}$ are not arbitrary: they arise naturally from the demand that the differential structure be compatible with the homothetic transformations introduced earlier. Specifically, the exterior derivative and codifferential must transform covariantly under the homothetic action $\Lambda$, preserving the structure of the deformed forms $\hat{\phi}=\Lambda\bar{\phi}$. This compatibility is encoded in the identities (29) which imply that homothetic transformations commute with the differential operators, effectively rendering the homothetic structure gauge covariant.
Physically, this suggests that the transformation group $\mathcal{H}$ plays a role analogous to a gauge group, albeit non-compact and infinite-dimensional. The field $\lambda(x)$ behaves like a scalar gauge parameter—more precisely, a logarithmic scale factor—whose differential $d\lambda$ enters the transformation of both $d$ and $\delta$ in a manner reminiscent of a connection 1-form in Weyl geometry. The antisymmetric contributions involving $d\lambda\wedge$ and $d\lambda\wedge*$ indicate that this is a kind of abelian connection with scale-coupled torsion.
@@
[…]
Figure 2: Diagrammatic representation of the generalized Hodge-de Rham complex with homothetic differential forms $\hat{\Omega}^{k}$, showing both the generalized exterior derivative $\hat{d}$ and codifferential $\hat{\delta}$. At each degree $k$, the kernel of the generalized Laplacian $\hat{\Delta}$ identifies the harmonic representatives $\mathcal{H}^{k}$ in the generalized Hodge decomposition $\mathcal{H}^{k}:=\ker\hat{\Delta}^{k}$ (${\mathcal{H}}^{k}$ not to be confused with the homothety group $\mathcal{H}$).
Cohomological Consequences
Importantly, the nilpotency of $\hat{d}$ allows us to define a generalized de Rham cohomology:
@@ 5 Homothetic Electromagnetism
The above homothetic Hodge-de Rham complex can be constructed for the electromagnetism, as a $U(1)$ gauge theory defined on the Minkowski spacetime manifold. It means embedding any $k$-form $\alpha^{k}$, where $0\leq k\leq 4$ in $\Omega^{k}(M)\oplus\Omega^{k}(M)$, where $\mathcal{H}$ can act on, making $\hat{\Omega}^{k}$. However, in order for the homothetic transformations to preserve the symmetries of the electromagnetism, they must act directly on the connections of the principal bundle, that is, the gauge potential fields $A=A_{\mu}dx^{\mu}$ (as the $\mathfrak{u}(1)$-valued 1-forms on $M$). Such a setup results in Fig. 3, in which we defined:
@@ 5.1 Sourceless Maxwell Equations
Then, $\hat{F}=\hat{d}\hat{A}$ is invariant under the homothetic gauge transformation:
$\boxed{\hat{A}\mapsto\hat{A}+\hat{d}\hat{\phi}}\qquad\text{where}\qquad\hat{\phi}=\Lambda\bar{\phi}=\Lambda\begin{pmatrix}\phi\\ \phi_{d}\end{pmatrix}$, (38)
and $\phi,\phi_{d}$ are arbitrary smooth functions on $M$. This is the homothetic equivalent of the classical gauge invariance of the Maxwell equations.
The first row of the homothetic Maxwell equations (36)-(37) yields:
@@
As one can see, some additional terms have appeared which represent the interaction of the electromagnetic fields with the offset fields. Apparently, for $d\lambda\equiv 0$, these additional terms disappear, and the classical Maxwell equations retrieve.
[…]
Figure 3: Homothetic Hodge-de Rham complex for electromagnetism (the lower part), and its relation with the classical complex (the upper part). The homothetic transformation genuinely occurs on the connection 1-forms, and it induces other transformations on the other $k\neq 1$ differential forms.
By translation of the above equations on the tangent bundle, the generalized Maxwell equations are obtained in the more conventional three-vector formulation:
@@
$\partial_{t}\tilde{E}-\nabla\times\tilde{B}+\partial_{t}\lambda(\tilde{E}-E_{d})-\nabla\lambda\times(\tilde{B}-B_{d})$ $=0,\qquad\text{generalized Ampère's Law};$ (42)
@@
$\nabla\cdot\tilde{B}=0,\quad\text{for }\nabla\cdot B_{d}$ $=0,\qquad\text{Gauss's Law for Magnetism}.$ (44)
This formulation offers a homothetically covariant generalization of electromagnetism. A distinguishing feature of this system is that it is sourceless. In fact, the emergence of penalty terms, with geometric origins, enable us to derive the dynamics of electromagnetic fields without the need for source terms. This issue is crucial in construction of a non-singular Model for a Point Electric Charge in $\S$ 8.2. However, for the sake of completeness, a non-homogeneous formulation is also provided in $\S$ 7.
6 The Homothetic Lorenz gauge, Wave, and Laplace Equations
We now turn to the derivation of the dynamical equations that arise from homothetically transformed electromagnetic potentials and scalar fields. These include a wave equation and a Laplace equation defined on the homothetically transformed geometry.
@@ 6.1 Homothetic Harmonicity of the Gauge Potential
Then, from the generalized Lorenz gauge (45), with:
@@
This yields a generalized conservation-like equation:
@@
In spacetime index notation:
@@ 6.2 Wave and Laplace Equations
7 Mathematics or Physics— HGT and the Spontaneous Symmetry Breaking
The principle of local scale invariance is a long-standing concept in theoretical physics. In Weyl’s original theory, fields transform multiplicatively. HGT proposes a different, affine-style transformation: $\phi\mapsto\phi_{d}+e^{-\lambda(x)}(\phi-\phi_{d})$. This introduces a fixed-point offset field, $\phi_{d}$, fundamentally altering the symmetry’s structure. This section argues that demanding invariance under this homothetic action is a form of spontaneous symmetry breaking (SSB). Meanwhile, the dynamics of $\lambda$ and the offset fields $\phi_{d}$ remain as the open questions.
First, it is important to recall two logically distinct types of scale transformations. In the classical Weyl approach [25], one considers local rescalings of the metric and fields as:
$g_{\mu\nu}\mapsto e^{2\sigma(x)}g_{\mu\nu},\qquad\phi\mapsto e^{W\sigma(x)}\phi,$ (48)
where $W$ is the Weyl weight of the field $\phi$. A non-Weyl-invariant model can be made invariant by introducing a compensating scalar field $\tau$ that transforms as $\tau\mapsto\tau-\sigma$. By dressing other fields with powers of $e^{\tau}$, one can achieve invariance under Weyl transformations. This is commonly interpreted as a spurionic or Stückelberg restoration of symmetry. Fixing a gauge (or assigning a vacuum expectation value to $\tau$) then breaks the symmetry spontaneously and realizes it nonlinearly. [^3] We ignore quantum anomaly issues here; the discussion is entirely classical.
In HGT, an analogous procedure applies, but the transformations act differently:
$g_{\mu\nu}\mapsto e^{-2\lambda(x)}g_{\mu\nu},\qquad\phi\mapsto\phi_{d}+e^{-\lambda(x)}(\phi-\phi_{d}),$ (49)
where $\lambda$ is the logarithm of the local scale factor, and $\phi_{d}$ is an offset field obeying the same equations as $\phi$ and transforming in the same representation. By defining $\delta\phi:=\phi-\phi_{d}$, the transformation simplifies to $\delta\phi\mapsto e^{-\lambda}\delta\phi$, meaning that $\delta\phi$ behaves like a conventional weight 1-1 field under local scalings. This encapsulates the essence of HGT transformations and allows the action to retain form-invariance under local homotheties.
Spontaneous symmetry breaking [11, 33] occurs when the underlying equations or action of a theory respect a symmetry, but the chosen vacuum state does not. In gauge theories, this typically manifests through the introduction of a scalar field acquiring a nonzero vacuum expectation value (VEV), thus selecting a specific vacuum and reducing the full symmetry to a residual subgroup. Importantly, SSB does not eliminate the symmetry but conceals it: physical configurations transform nontrivially under the broken generators. In classical field theory, this principle underlies both the Higgs mechanism and geometric symmetry reductions. Within the context of HGT, SSB becomes the bridge between mathematical symmetry and emergent physical structure, as is exemplified by the generalized Maxwell equations in the sequel.
In $\S$ 5.1, we derived the source-free generalized Maxwell equations directly from the homothetic Hodge–de Rham theory. This geometry-based approach establishes the mathematical foundation of the equations. Regardless of their physical interpretation, the homothetic gauge freedom and the derived equations are of interest in their own right, with potential applications to the theory of partial differential equations and to computational physics. Here we shall derive the same equations, in the presence of the current density $J\neq 0$, from a process that closely reflects spontaneous symmetry breaking.
The starting point is an action that is not manifestly scale-invariant due to explicit dependence on a metric and field configurations. By introducing the homothetic variables $(\phi,\phi_{d})$ and employing the transformation rule in Eq. (49), the action is modified in such a way that symmetry is effectively restored at the level of its form. This mirrors the logic of Stückelberg restoration: one modifies the theory so that it is invariant under a larger symmetry group, and the breaking is realized through a fixed gauge or background choice.
This viewpoint implies that HGT treats the effect of local scale transformation as a spontaneous symmetry breaking mechanism. However, unlike traditional gauge theories with Higgs-like scalars, the dynamical origins and physical interpretations of the fields $\lambda$ and $\phi_{d}$ remain unresolved. The field $\lambda(x)$ governs local rescalings but is not yet derived from a fundamental action. Likewise, the offset field $\phi_{d}$ is introduced to enable covariant transformations, but its dynamics and role in observable quantities have not been clarified. These gaps must be addressed if HGT is to evolve into a physical rather than purely formal framework.
7.1 Example: The Generalized Maxwell Equations and Scale-symmetry Restoration
For our derivations, we recall the following relations (from $\S$ 4.2):
  1. 1.
    The block differential on $k$-forms,
    $\hat{d}\begin{pmatrix}\alpha\\ \beta\end{pmatrix}\ :=\ \begin{pmatrix}d\alpha+b\wedge\alpha-b\wedge\beta\\[2.0pt] d\beta\end{pmatrix},\qquad b:=d\lambda,$ (50)
    and the adjoint $\hat{\delta}:=-\hat{*}\hat{d}\hat{*}$ with respect to $\tilde{\eta}$, where $\hat{*}$ is the Hodge operator on the block bundle built from $_{\tilde{\eta}}$ on each component.
  2. 2.
    The block field-strength $\hat{F}:=\hat{d}\,\hat{A}$ for a one-form doublet $\hat{A}:=\big{(}\tilde{A},A_{d}\big{)}^{\mathsf{T}}$, with upper component
    $\tilde{F}:=d\tilde{A}+b\wedge(\tilde{A}-A_{d}),\qquad F_{d}:=dA_{d}.$ (51)
Start from the standard Maxwell system on $(M,\eta)$ coupled to a conserved current $J$,
$S_{\rm Max}[A;J]\ =\ -\frac{1}{2}\int F\wedge*_{\eta}F\ +\ \int A\wedge*_{\eta}J,\qquad F:=dA,\quad d*_{\eta}J=0,$ (52)
which in $d=4$ is conformally invariant in the sourceless sector but, in the presence of sources and a fixed choice of scale, need not be Weyl-invariant in the sense of (48). To restore local scale symmetry homothetically, introduce the compensator $\lambda$, the offset potential $A_{d}$, and the doublet $\hat{A}=(\tilde{A},A_{d})^{\mathsf{T}}$ with
$\tilde{A}\ :=\ A_{d}+e^{-\lambda}(A-A_{d}),\qquad\text{so that}\qquad\tilde{A}-A_{d}=e^{-\lambda}(A-A_{d}).$ (53)
We then define the HGT Maxwell action by
$S_{\rm HGT}[\,\hat{A};\lambda;\hat{J}\,]\ :=\ -\frac{1}{2}\int\big{\langle}\hat{F},\hat{F}\big{\rangle}_{\tilde{\eta}}\ +\ \int\big{\langle}\hat{A},\hat{J}\big{\rangle}_{\tilde{\eta}},\qquad\hat{F}:=\hat{d}\,\hat{A}$, (54)
where $\hat{J}:=(\tilde{J},J_{d})^{\mathsf{T}}$ is a source doublet (the top slot couples to the physical sector; the bottom slot is available for regularizations/defects and can be set to zero in the simplest applications). By construction, $S_{\rm HGT}$ is exactly invariant under local homotheties (49) with $b=d\lambda$, since all derivatives are the Weyl-covariant ones and the pairing uses $\tilde{\eta}$.
Varying $\hat{A}$ yields the generalized Maxwell equations
$\boxed{\quad\hat{d}\,\hat{F}\ =\ 0,\qquad\hat{\delta}\,\hat{F}\ =\ \hat{J},\quad}$ (55)
i.e. the Weyl-covariant Bianchi identity and source equation on the block bundle. In a Weyl-covariant Lorenz gauge,
$\hat{\delta}\,\hat{A}=0\quad\Longleftrightarrow\quad\tilde{\nabla}_{\mu}\tilde{A}^{\mu}\;-\;b_{\mu}\big{(}\tilde{A}^{\mu}-A_{d}^{\mu}\big{)}\ =\ 0,$ (56)
the equations reduce to a block wave equation
$\hat{\Delta}\,\hat{A}\ =\ \hat{J},\qquad\hat{\Delta}:=\hat{d}\,\hat{\delta}+\hat{\delta}\,\hat{d}$, (57)
with $\hat{\Delta}$ the HGT Laplace–de Rham operator built from $\tilde{\eta}$ and $b=d\lambda$. Writing out only the physical (upper) slot, one obtains the compact form
$\delta_{\tilde{\eta}}\!\left[d\tilde{A}+b\wedge(\tilde{A}-A_{d})\right]\ =\ \tilde{J},\qquad d\left[d\tilde{A}+b\wedge(\tilde{A}-A_{d})\right]\ =\ 0,$ (58)
where $\delta_{\tilde{\eta}}$ is the codifferential with respect to $\tilde{\eta}=e^{-2\lambda}\eta$. When $A_{d}$ is chosen as a pure gauge or fixed reference (e.g. $A_{d}\equiv 0$) these reduce to
$\delta_{\tilde{\eta}}\!\left(d\tilde{A}+b\wedge\tilde{A}\right)=\tilde{J},\qquad d\left(d\tilde{A}+b\wedge\tilde{A}\right)=0,$ (59)
which are the Maxwell equations in an integrable Weyl geometry with scale connection $b=d\lambda$. Equations (56)–(59) make manifest that the only effect of the restoration is to (i) replace ordinary by Weyl-covariant differentials and (ii) introduce a mixing with the offset field $A_{d}$ through $\tilde{A}-A_{d}$, i.e. through the homogeneously transforming difference.
8 Applications
From here on, some applications of the theory will be discussed. Two related mathematically and physically important problems will be tackled in the framework of the homothetic gauge theory, and some new results will be provided and discussed.
8.1 Cauchy, Dirichlet and Neumann Boundary Conditions for the Laplace Equation
According to the left part of Fig. 4, let $S\subset\mathbb{R}^{d}$ be a closed, smooth, orientable, embedded (d1)(d-1)-dimensional submanifold, where $d\in\{1,2,3\}$. By the Jordan–Brouwer separation theorem, $\mathbb{R}^{d}\setminus S$ has exactly two connected components: one bounded (denote it by $\Omega_{i}$) and one unbounded (denote it by $\Omega_{o}$). The closures $\overline{\Omega_{i}}$ and $\overline{\Omega_{o}}$ are smooth $d$-dimensional manifolds with boundary, sharing the boundary $\partial\overline{\Omega_{i}}=\partial\overline{\Omega_{o}}=S$. Then, we consider $S_{i}$ and $S_{o}$ as the interior and exterior sides of $S$, respectively.
Note 1
From here on, we provide a general formulation, which is valid equally for both the interior manifold $\overline{\Omega_{i}}$ and the exterior manifold $\overline{\Omega_{o}}$. Therefore, we avoided repeatedly using of indices ”ii” and ”oo”.